Find two integers whose product is 104 such that one of the integers is 3 less than twice the other integer. Thank you for showing me how to set this problem up.
Answer by edjones(8007)

x-3=2y
x=2y+3
.
xy=104
y(2y+3)=104
2y^2+3y-104=0
2*-104=-208. The 2 factors of -208 whose sum is 3 are -13 and 16.
2y^2+16y-13y-104=0
2y(y+8)-13(y+8)=0 Factoring by grouping.
(2y-13)(y+8)=0
.
y=-8
2y=13, y=13/2 not an integer. Did the question ask for two numbers or integers?
.
Ed